Data Set: 100,96,94,94,92,88,87,86,86,86,82,70
1. What is the z score for a score of 100?
A. 1.488
B. 0.974
C. 1.503
D. 1.000

2. What is the z score for a score of 94?
A. 0.717
B. 0.608
C. 0.689
D. 0.460

3. What is the z score for a score of 86?
A. -0.359
B. -0.310
C. 0.460
D. -0.615

4. What is the mean of the z scores?
A. 1
B. 0.321
C. 0
D. 0.50




sidenote: i have tried putting the numbers in my calc and none of my answers are these options provided for my homework. PLEASE HELP ME THANK YOU   
Answer #1

We have given the data.

Data Set: 100,96,94,94,92,88,87,86,86,86,82,70

From the given data we get -

Mean x bar = 88.4167 and standard deviation s = 7.7864

Z score for sample value x -

Z score =( x - mean ) / s

1. For x = 100

Z score = (100 - 88.4167) / 7.7864 = 1.488

Hence   A. 1.488 is the correct answer.

2. For x = 94

Z score = (94 - 88.4167) / 7.7864 = 0.717

Hence   A. 0.717 is the correct answer.

3. For x = 86

Z score = (86 - 88.4167) / 7.7864 = -0.310

Hence   B. -0.310 is the correct answer.

4. We know that for standard normal distribution the mean is 0 and standard deviation is 1.

Hence mean of the z scores is 0

Hence C. 0 is the correct answer.

Hope this will help you. Thank you :)